# Generalized Recurrence Criteria for Classes of Open Quantum Walks

### v1

@@ Abstract
In this paper, we study the recurrence of open quantum walks (OQWs) induced by finite-dimensional coins $(L,B,R)$. The focus is on homogeneous OQWs with a set of vertices $\mathbb{Z}$, the set of integers. We present three distinct recurrence criteria, each adapted to different types of coins. The first criterion was developed for a class of Lazy OQWs in any finite dimension, where the presented criterion is associated with an auxiliary map and its only invariant state, resulting in the first recurrence criterion for Lazy OQWs. The second one is restricted to Lazy OQWs of dimension 2, where we provide a complete characterization of the recurrence for this lower dimension. Finally, we present a general criterion for finite-dimensional coins in the non-lazy case $(B=0)$, which generalizes many of the previously known results. This new criterion holds for irreducible and reducible OQWs through a decomposition of the Hilbert space where our quantum states act.
@@
Open Quantum Walks, Recurrence, Lazy walks, Coin, General Criterion
@@ 1 Introduction
Open Quantum Walks (OQWs) represent a natural extension of classical Markov chains through open quantum systems. Originally proposed by [2], OQWs are models that capture the behavior of open quantum systems, in which the measurement of the particle position after each time step leads to a classical Markov chain when it is joined with the space of particle states.
In this context, we consider homogeneous OQWs defined on the set of integers, that is, the graph of this type of random walk is the set of vertices $V=\mathbb{Z}$, where the linear operators $R$ and $L$ play the roles of transitions between the vertices of the quantum walk, which are constant and each jump only occurs from one vertex to its nearest neighbors. Homogeneous OQWs in the Lattice are explored and detailed in [6].
Recurrence is an important statistic of any random walk, emerging as a central question of interest in the quantum context, and can be directly connected with the classical recurrence criteria for random walks in $\mathbb{Z}$. For example, for homogeneous classical random walks on $\mathbb{Z}$, it is well known that a walk is recurrent if and only if the transition parameters $r$ (rate of jumping to right) and $l$ (rate of jumping to left) are equal (or $r=1/2=l$ in the case $b=0$) [11]. This classical random walk has its graph illustrated in Figure 1.
@@
Figure 1: Classical Lazy Random Walk on $\mathbb{Z}$.
With this description, it is natural to seek criteria to determine under what conditions a homogeneous OQW is recurrent given $R$ and $L$. This problem has been addressed in the literature with the development of specific criteria, which are essential to understanding the dynamics and long-term behavior of these systems.
A Criterion for PQ Matrices can be found in [10]. Also, in that work the authors analyze the positive recurrence of this type of walk. From the analysis of the spectrum of $L^{*}L$ and $R^{*}R,$ it was showed in [7] that $R$ and $L$ having only singular values equal to $1/\sqrt{2}$ is a sufficient condition to guarantee that those OQWs are recurrent. More recently, it was obtained general criteria for recurrence of non-Lazy OQWs in dimension 2 and for irreducible finite-dimensional OQWs [9].
In this article, we present a general criterion for the recurrence of homogeneous OQWs, based on the transition operators $R$ and $L$ when acting in finite-dimensional vector spaces, that is, we allow the OQW to be reducible in any finite dimension. Furthermore, we present the first recurrence criterion for lazy OQWs, using results obtained from recent studies on Central Limit Theorems for this kind of OQWs.
A first Central Limit Theorem for OQWs was presented by [1]. Here we will use more recent results. These works presented a Central Limit Theorem for Lazy OQWs [8] and one for any homogeneous OQW in Lattice when the degrees of freedom are finite [5].
This work is structured as follows. In Section 2 we define the open quantum walks induced by a coin and their transition probabilities. Section 3 is devoted to presenting a recurrence criterion for a class of Lazy OQWs, and one restricted to Lazy OQWs of dimension 2. In Section 4 we focus on the case where the coin is non-lazy and we present a general recurrence criterion for finite-dimensional coins. Section 5 illustrates the results with comprehensive examples.
@@ 2 General Settings
Let us consider a set of vertices $V$, which might be finite or countable infinite. Let $\mathcal{K}=\mathbb{C}^{V}$ be the state space of a quantum system with as many degrees of freedom as the number of vertices. In the case where $V$ is infinite countable, we assume $\mathcal{K}$ to be any separable Hilbert space with an orthonormal basis indexed by $V$, denote by $\left\{\lvert i\rangle\right\}_{i\in V}$. To emphasize the position of the particle of our quantum walk, we consider a quantum system described by the separable complex Hilbert space
$\mathcal{H}=\bigoplus_{i\in V}\mathfrak{h}_{i},$
where the $\mathfrak{h}_{i}$ are separable Hilbert spaces. The label $i\in V$ represents the position of the particle and, when the particle is located at the vertex $i\in V$ , its internal state is encoded in the space $\mathfrak{h}_{i},$ as it will be detailed below.
We consider the set of density operators over $\mathcal{H},$
$\mathcal{D}(\mathcal{H})=\{\rho\in\mathcal{H}:\rho\geq 0,\mathrm{Tr}(\rho)=1\},$
@@
$\mathcal{S}(\mathcal{H}\otimes\mathcal{K})=\left\{\tau\in\mathcal{I}(\mathcal{H}\otimes\mathcal{K}):\textstyle\tau=\displaystyle\sum_{j\in\mathbb{Z}}\rho(j)\otimes\lvert j\rangle\langle j\rvert\right\},$
the set of density operators acting on $\mathcal{H}\otimes\mathcal{K}.$ Fixed a set of vertices $V,$ an open quantum random walk (OQW) on $V$ is a completely positive map
$\Phi:\mathcal{H}\otimes\mathcal{K}\mapsto\mathcal{H}\otimes\mathcal{K},$
given by
$\Phi(\tau)=\sum_{i\in V}\Bigg{(}\sum_{j\in V}B_{i}^{j}\rho_{j}B_{i}^{j^{*}}\Bigg{)}\otimes|i\rangle\langle i|,\;\;\;\tau\in\mathcal{S}(\mathcal{H}\otimes\mathcal{K}),$
where $B_{i}^{j}$ are bounded operators on $\mathcal{H}$ corresponding to the transition effect from site $j$ to $i$, and
$\sum_{i\in V}B_{i}^{j*}B_{i}^{j}=I,\;\;\;j\in V,$
which guarantees that $\mathrm{Tr}(\Phi(\tau))=\mathrm{Tr}(\tau)$, for every $\tau\in\mathcal{D}(\mathcal{H})$, meaning that $\Phi$ is trace-preserving, thus $\Phi$ preserves the set $\mathcal{S}(\mathcal{H}\otimes\mathcal{K}),$ i.e., $\Phi(\tau)\in\mathcal{S}(\mathcal{H}\otimes\mathcal{K})\;\forall\tau\in\mathcal{S}(\mathcal{H}\otimes\mathcal{K}).$
In this work, we focus on the OQW model presented in [2], where the set of vertices is $V=\mathbb{Z}.$ For simplicity, we will say integer line, or simply the line, to refer to the set $\mathbb{Z}$ of integer numbers. An OQW on $\mathbb{Z}$ concerns the evolution of density operators of the form
$\tau=\sum_{j\in\mathbb{Z}}\rho(j)\otimes\lvert j\rangle\langle j\rvert\;\in\;\mathcal{H}\otimes\mathbb{C}^{\mathbb{Z}},$
where the $V=\mathbb{Z}$ represents the position of the walker and $\mathfrak{h}_{i}$ is the Hilbert space describing the internal degrees of freedom for site $\lvert i\rangle$, which we call the coin space and will be denoted by $\mathfrak{h}_{i}=\mathbb{C}^{k}$. Finally, we consider homogeneous lazy open quantum walks: at each step either the particle on site $\lvert j\rangle$ moves exactly one position to some of its two nearest-neighbours, in the left ($\lvert j-1\rangle$) or in the right ($\lvert j+1\rangle$), or it remains at same site ($\lvert j\rangle$), and the coin is constant, i.e., the same transition rule is applied at each site. Therefore, the linear operator $\Phi$ can be written by
$\Phi(\tau)=\sum_{i\in\mathbb{Z}}\sum_{j=i-1}^{i+1}M_{i}^{j}\tau(M_{i}^{j})^{*},$
where
$M_{i}^{i-1}=L\otimes\lvert i-1\rangle\langle i\rvert,\quad M_{i}^{i}=B\otimes\lvert i\rangle\langle i\rvert,\quad M_{i}^{i+1}=R\otimes\lvert i+1\rangle\langle i\rvert.$
The linear operators $L,B,R\in\mathbb{M}_{k}({\mathbb{C}})$ are matrices such that $L^{*}L+B^{*}B+R^{*}R=I_{k},$ and simplified by
$\Phi(\tau)=\sum_{i\in\mathbb{Z}}\Big{(}L\rho(i+1)L^{*}+B\rho(i)B^{*}+R\rho(i-1)R^{*}\Big{)}\otimes|i\rangle\langle i|,\;\;\;\tau\in\mathcal{S}(\mathcal{H}\otimes\mathcal{K}).$ (1)
@@
If $B\neq 0,$ the quantum walk exhibits a lazy behavior, meaning that in the context of quantum walks there is a non-zero probability for the walker to stay in the same vertex after a step. We refer to the OQW above as an OQW induced by the coin $(L,B,R)$ in which the walk is inherently lazy. This lazy behavior introduces distinct properties compared to its non-lazy counterpart. Notably, the presence of the matrix $B$ influences the walk’s propensity to remain in its current vertex, thereby affecting the overall dynamics. If $B=0,$ then $\Phi$ will be called an OQW induced by a coin $(L,R)$ or an non-lazy OQW induced by a coin $(L,R).$
The dynamic described above is represented by the graph in Figure 2.
@@
The OQW induced by a coin $(L,B,R)$ is the homogeneous, nearest-neighbor OQW on the integer line with the left and right transitions given by $L$ and $R$, respectively, and $B$ represents the rate of no jump at some instant. In this sense, the quantum trajectory of this kind of OQW on the line, induced by a coin $(L,B,R)$ and starting from a state $\tau$ of the form $\tau=\sum_{i\in\mathbb{Z}}\rho(i)\otimes|i\rangle\langle i|,$ is any path generated by the Markov chain $(X_{n},\tau_{n})_{n\geq 0}$, where $X_{n}$ denotes the position of the particle at time $n$ and $\tau_{n}$ its internal degree. The transition probabilities are given by
$\begin{split}\mathbb{P}\left(\,(X_{n+1},\tau_{n+1})=\left(i-1,\frac{L\sigma L^{*}}{\mathrm{Tr}(L\sigma L^{*})}\right)\;\Bigg{|}\;(x_{n},\rho_{n})=(i,\sigma)\,\right)&=\mathrm{Tr}(L\sigma L^{*}),\\ \mathbb{P}\left(\,(X_{n+1},\tau_{n+1})=\left(i,\frac{B\sigma B^{*}}{\mathrm{Tr}(B\sigma B^{*})}\right)\;\Bigg{|}\;(x_{n},\rho_{n})=(i,\sigma)\,\right)&=\mathrm{Tr}(B\sigma B^{*}),\\ \mathbb{P}\left(\,(X_{n+1},\tau_{n+1})=\left(i+1,\frac{R\sigma R^{*}}{\mathrm{Tr}(R\sigma R^{*})}\right)\;\Bigg{|}\;(x_{n},\rho_{n})=(i,\sigma)\,\right)&=\mathrm{Tr}(R\sigma R^{*}).\end{split}$
for every $i\in\mathbb{Z}$, $\sigma\in\mathcal{S}(\mathfrak{h}_{i})$, and initial law
$\mathbb{P}\left(\,(X_{0},\rho_{0})=\left(i,\frac{\rho(i)}{\mathrm{Tr}\,\rho(i)}\right)\;\right)=\mathrm{Tr}(\rho(i)),$
and all other transition probabilities are null.
@@ 3.1 Faithfulness of Density Operators
Density operators play a fundamental role in open quantum random walks, providing a statistical description of this type of quantum system. Faithful and non-faithful density operators are two categories that arise in the study of quantum states, being characterized by their positive definiteness and positive-semidefiniteness, respectively.
Faithful density operators are defined as being positive definite. This means that the eigenvalues of a faithful density operator are strictly positive and it is invertible. In essence, faithfulness ensures a faithful representation of the quantum system, capturing the full richness of its physical properties.
On the other hand, non-faithful density operators are positive-semidefinite, thus they have $\lambda=0$ as one of their eigenvalues. Thereby, that there may be ambiguity in the quantum state description, as multiple states can correspond to the same non-faithful density operator. The positivity-semidefiniteness criterion relaxes the strictness of positive definiteness, permitting a broader range of quantum state representations.
An important subset of the set of non-faithful density operators is composed of the pure density operators. A pure density is a density $\rho$ that is a projection of rank 1, i.e., there exists some vector $\lvert u\rangle\in\mathbb{C}^{k}$ such that $\rho=\lvert u\rangle\langle u\rvert$.
@@
$\Phi^{n}(\rho\otimes\lvert i\rangle\langle i\rvert)=\sum_{s\in V}\rho_{n}(s)\otimes\lvert s\rangle\langle s\rvert.$
@@
denote the probability of being at site $j$ at step $n,$ given that we started at site $i,$ with initial density $\rho$ concentrated at $i.$ Therefore, the dynamics starts with a density operator $\rho$ concentrated at some vertex $\lvert i\rangle,$ takes the evolution up to time ($n$) through the linear operator $\Phi^{n},$ producing a new density operator
@@
Fixed some $i\in V,$ we denote the set of density operators acting on $\mathfrak{h}_{i}$ by
$\mathcal{S}(\mathfrak{h}_{i})=\left\{\rho:\mathfrak{h}_{i}\rightarrow\mathfrak{h}_{i}:\;\mathrm{Tr}(\rho)=1,\;\rho=\rho^{*};\;\rho\geq 0\right\}.$
Consider an OQW on some set of vertices $V$, let $\lvert i\rangle\in V$ and $\rho\in\mathcal{S}(\mathfrak{h}_{i})$. We say that a vertex $\lvert i\rangle$ is
@@
    $\sum_{n=0}^{\infty}p_{ii;\rho}(n)=\infty.$
    Otherwise, $i$ is said to be $\rho$-transient;
@@
    recurrent, if $i$ is $\rho-$recurrent for all $\rho\in\mathcal{S}(\mathfrak{h}_{i});$
@@
    transient, if $i$ is $\rho-$transient for all $\rho\in\rho\in\mathcal{S}(\mathfrak{h}_{i}).$
Concerning all vertices, an OQW is called: recurrent if every vertex is recurrent; transient if every vertex is transient.
The following Proposition describes $\rho$-recurrence in a concentrated vertex $\lvert i\rangle\in V$ when assumed that $\lvert i\rangle$ is recurrent for some density and $\mathfrak{h}_{i}$ has a finite internal degree of freedom. An immediate consequence will also be given for the case where $\lvert i\rangle$ is $\sigma$-transient for some non-faithful density.
@@
Consider an OQW on some set of vertices $V,$ let $i\in V$ and assume that $\mathfrak{h}_{i}$ is finite-dimensional. If $\lvert i\rangle$ is $\tau$-recurrent for some $\tau\in\mathcal{S}(\mathfrak{h}_{i}),$ then
@@
    vertex $\lvert i\rangle$ is $\rho$-recurrent for all faithful $\rho\in\mathcal{S}(\mathfrak{h}_{i})$;
@@
    there exists a non-faithful density operator $\tilde{\rho}$ such that $\lvert i\rangle$ is $\tilde{\rho}$-recurrent;
  3. 3.
    there exists a pure density operator $\lvert w\rangle\langle w\rvert$ such that the OQW is $\lvert w\rangle\langle w\rvert$-recurrent, where $\lvert w\rangle$ is an unitary vector of $\mathfrak{h}_{i}.$

### v2

# Recurrence Criteria for Reducible Homogeneous Open Quantum Walks on the Line

@@ Abstract
In this paper, we study the recurrence of homogeneous Open Quantum Walks (HOQWs) induced by finite-dimensional coins $(L,B,R)$. We present three distinct recurrence criteria, each adapted to different types of coins. The first criterion applies to coins whose auxiliary map has a unique invariant state, resulting in the first recurrence criterion for Lazy HOQWs. The second one applies to Lazy HOQWs of dimension 2, where we provide a complete characterization of the recurrence for this low-dimensional case. Finally, we present a general criterion for finite-dimensional coins in the non-lazy case $(B=0)$, which generalizes many of the previously known results. This new criterion holds for irreducible and reducible HOQWs.
@@
Homogeneous Open Quantum Walks, Recurrence, Lazy Walks, Coin, Reducibility
@@ 1 Introduction
Open Quantum Walks (OQWs) naturally extend classical Markov chains within the framework of open quantum systems. Originally proposed by [2], OQWs describe open quantum systems in which the repeated measurement of the particle’s position, together with the space of particle states, gives rise to a Markov chain.
In this context, we consider homogeneous OQWs (HOQWs) defined on the set of integers, that is, the graph of this type of random walk is the set of vertices $V=\mathbb{Z}$, where 3 linear operators $L,B,R$ play the roles of transitions between the vertices of the quantum walk, which are constant and each jump only occurs from one vertex to its nearest neighbors, or to itself. HOQWs on more general lattices are explored and detailed in [7].
Recurrence is an important probabilistic property of any random walk, often characterized by the expected number of returns to the initial vertex, emerging as a central question of interest in the quantum context, and can be directly connected with the classical recurrence criteria for random walks in $\mathbb{Z}$. For example, consider the classical homogeneous random walk on $\mathbb{Z}$ with graph given by Figure 1. It is well known that this lazy random walk with $r,l\geq 0,\;b>0,\;l+b+r=1,$ is recurrent if and only if the transition parameters $r$ (rate of jumping to right) and $l$ (rate of jumping to left) are equal. Also, if $b=0,$ then the walk is non-lazy and it is recurrent if and only if $r=1/2=l$ [14].
@@
Figure 1: Classical Homogeneous Lazy Random Walk on $\mathbb{Z}$.
In the context of discrete-time Markov chains, a state is called recurrent if, once the chain enters that state, the probability of returning to it is 1. This is equivalent to saying that the average number of returns to that state is infinite, indicating that the chain will revisit the state indefinitely [14].
With this description, it is natural to seek criteria to determine under what conditions a HOQW is recurrent given $L,B,R.$ This problem has been addressed in the literature with the development of specific criteria, which are essential to understanding the dynamics and long-term behavior of these systems. See for instance [8, 13, 12]. These works define recurrence in terms of probability of return, where a state is said to be recurrent if the probability of the walk returning to it is equal to 1. However, it has been shown that for OQWs, the two definitions of recurrence based on returning with probability 1 is not equivalent to the one based on the average number of returns being infinite [3]. This property highlights a distinct characteristic of OQWs compared to classical Markov chains.
Once this difference is understood, we discuss the recurrence criteria for OQWs, adopting the definition proposed in [3]. Specifically, we analyze the recurrence of OQWs based on the average number of returns.
Furthermore, recurrence criteria concerning the mean number of returns for higher-order HOQWs have been presented in [11]. The authors of that work extend the analysis of recurrence to a more general class of open quantum walks, such as irreducible coins.
In this article, we present more general criteria for the recurrence of lazy and non-lazy HOQWs, based on the transition operators $L,B,R$ when acting on finite-dimensional vector spaces, that is, we allow the HOQW to be reducible. Furthermore, we present the first recurrence criterion for lazy HOQWs, using results obtained from recent studies on Central Limit Theorems for this kind of HOQWs.
A first Central Limit Theorem (CLT) for Open Quantum Walks (OQWs) was established in [1]. In this work, we rely on more recent developments, including the CLT for Lazy OQWs presented in [10], as well as the general CLT for homogeneous OQWs (HOQWs) on lattices with finite internal degrees of freedom given in [6].
This work is structured as follows. In Section 2 we define the open quantum walks induced by a coin and their transition probabilities. Section 3 is devoted to presenting a recurrence criterion for a class of Lazy HOQWs, and one restricted to Lazy HOQWs of dimension 2. In Section 4 we focus on the case where the coin is non-lazy and we present a general recurrence criterion for finite-dimensional coins. Section 5 illustrates the results with comprehensive examples.
@@ 2 General Settings
Let us consider the set of vertices $\mathbb{Z},$ and $\mathcal{K}=\mathbb{C}^{\mathbb{Z}}$ be the state space of a quantum system with as many degrees of freedom as the number of vertices. To emphasize the position of the particle of our quantum walk, we consider a quantum system described by the separable complex Hilbert space
$\mathcal{H}=\bigoplus_{i\in\mathbb{Z}}\mathfrak{h}_{i},$
where the $\mathfrak{h}_{i}$ are separable Hilbert spaces. The label $i\in\mathbb{Z}$ represents the position of the particle and, when the particle is located at the vertex $i\in\mathbb{Z}$, its internal state is encoded in the space $\mathfrak{h}_{i},$ as it will be detailed below.
We consider the set of density operators over a finite Hilbert space $\mathcal{W},$
$\mathcal{D}(\mathcal{W})=\{\rho:\mathcal{W}\rightarrow\mathcal{W}:\;\rho\geq 0,\;\mathrm{Tr}(\rho)=1\},$
@@
$\mathcal{S}(\mathcal{H}\otimes\mathcal{K})=\left\{\tau\in\mathcal{I}(\mathcal{H}\otimes\mathcal{K}):\textstyle\tau=\displaystyle\sum_{j\in\mathbb{Z}}\rho(j)\otimes\lvert j\rangle\langle j\rvert,\quad\rho(j)\geq 0,\quad\sum_{j\in\mathbb{Z}}\mathrm{Tr}\left(\rho(j)\right)=1\right\},$
the set of density operators acting on $\mathcal{H}\otimes\mathcal{K}.$ We are interested in HOQWs of the form
$\Phi(\tau)=\sum_{i\in\mathbb{Z}}\Big{(}L\rho(i+1)L^{*}+B\rho(i)B^{*}+R\rho(i-1)R^{*}\Big{)}\otimes|i\rangle\langle i|,\;\;\;\tau\in\mathcal{S}(\mathcal{H}\otimes\mathcal{K}),$ (1)
where $L,B,R\in\mathbb{M}_{d}({\mathbb{C}})$ satisfy $L^{*}L+B^{*}B+R^{*}R=I_{d}.$
@@
If $B\neq 0,$ the quantum walk exhibits a lazy behavior, meaning that in the context of quantum walks there is a non-zero probability for the walker to stay in the same vertex after a step. We refer to the HOQW above as a HOQW induced by the coin $(L,B,R)$ in which the walk is inherently lazy.
Therefore, this coin represents a homogeneous, nearest-neighbor OQW on the integer line with the left and right transitions given by $L$ and $R$, respectively, and $B$ represents the rate of no jump at some instant.
This lazy behavior introduces distinct properties compared to its non-lazy counterpart. Notably, the presence of the matrix $B$ influences the walk’s propensity to remain in its current vertex, thereby affecting the overall dynamics. If $B=0,$ then $\Phi$ will be called a HOQW induced by a coin $(L,R)$ or a non-lazy HOQW induced by a coin $(L,R).$ The dynamic described above is represented by the graph in Figure 2.
@@
The quantum trajectory of this kind of OQW on the line, induced by a coin $(L,B,R)$ and starting from a state $\tau$ of the form $\tau=\sum_{i\in\mathbb{Z}}\tau_{i}\otimes|i\rangle\langle i|,$ is any path generated by the Markov chain $(X_{n},\tau_{n})_{n\geq 0}$, where $X_{n}$ denotes the position of the particle at time $n$ and $\tau_{n}$ its internal degree. The transition probabilities are given by
$\begin{split}\mathbb{P}\left(\,(X_{n+1},\tau_{n+1})=\left(i-1,\frac{L\sigma L^{*}}{\mathrm{Tr}(L\sigma L^{*})}\right)\;\Bigg{|}\;(x_{n},\tau_{n})=(i,\sigma)\,\right)&=\mathrm{Tr}(L\sigma L^{*}),\\ \mathbb{P}\left(\,(X_{n+1},\tau_{n+1})=\left(i,\frac{B\sigma B^{*}}{\mathrm{Tr}(B\sigma B^{*})}\right)\;\Bigg{|}\;(x_{n},\tau_{n})=(i,\sigma)\,\right)&=\mathrm{Tr}(B\sigma B^{*}),\\ \mathbb{P}\left(\,(X_{n+1},\tau_{n+1})=\left(i+1,\frac{R\sigma R^{*}}{\mathrm{Tr}(R\sigma R^{*})}\right)\;\Bigg{|}\;(x_{n},\tau_{n})=(i,\sigma)\,\right)&=\mathrm{Tr}(R\sigma R^{*}),\end{split}$
for every $i\in\mathbb{Z}$, $\sigma\in\mathcal{D}(\mathfrak{h})$, and initial law
$\mathbb{P}\left(\,(X_{0},\tau_{0})=\left(i,\frac{\tau_{i}}{\mathrm{Tr}\,\tau_{i}}\right)\;\right)=\mathrm{Tr}(\tau_{i}),$
and all other transition probabilities are null. Therefore, at time $n$, if the system is located at site $\lvert i\rangle$ with an initial density $\rho\otimes\lvert i\rangle\langle i\rvert$, then at the next time step $n+1$, one of the following occurs:
  • The particle moves to site $\lvert i-1\rangle$ (left) with probability $\operatorname{Tr}(L_{i}\rho L_{i}^{*})$, and the density evolves to
    $\frac{L_{i}\rho L_{i}^{*}}{\operatorname{Tr}(L_{i}\rho L_{i}^{*})}\otimes\lvert i-1\rangle\langle i-1\rvert;$
  • The particle remains at site $\lvert i\rangle$ with probability $\operatorname{Tr}(B_{i}\rho B_{i}^{*})$, and the new density matrix is
    $\frac{B_{i}\rho B_{i}^{*}}{\operatorname{Tr}(B_{i}\rho B_{i}^{*})}\otimes\lvert i\rangle\langle i\rvert;$
  • The particle moves to site $\lvert i+1\rangle$ (right) with probability $\operatorname{Tr}(R_{i}\rho R_{i}^{*})$, leading to the new density matrix
    $\frac{R_{i}\rho R_{i}^{*}}{\operatorname{Tr}(R_{i}\rho R_{i}^{*})}\otimes\lvert i+1\rangle\langle i+1\rvert.$
@@ 3.1 Faithfulness of Density Operators
Density operators play a fundamental role in open quantum systems. In this work, they provide statistical descriptions of HOQWs. Faithful and non-faithful density operators are characterized by their positive definiteness and positive semi-definiteness, respectively, when they evolve in finite-dimensional systems.
Faithful density operators are defined as being positive definite. This means that the eigenvalues of a faithful density operator are strictly positive, and the operator is invertible. Essentially, faithfulness ensures a complete representation of the quantum system, capturing the full richness of its physical properties.
On the other hand, non-faithful density operators are positive semi-definite, thus they have $\lambda=0$ as one of their eigenvalues. The presence of zero eigenvalues implies that the non-faithful density operator is not invertible and may represent specific mixed states or situations where information loss occurs. The positive semi-definiteness criterion relaxes the strictness of positive definiteness, allowing for a broader range of quantum state representations, while still maintaining the one-to-one correspondence between quantum states and density operators.
An important subset of the set of non-faithful density operators is composed of the pure density operators. A pure density is a density $\rho$ that is a projection of rank 1, i.e., there exists some vector $\lvert u\rangle\in\mathbb{C}^{d}$ such that $\rho=\lvert u\rangle\langle u\rvert$.
@@
$\Phi^{n}(\rho\otimes\lvert i\rangle\langle i\rvert)=\sum_{s\in\mathbb{Z}}\rho_{n}(s)\otimes\lvert s\rangle\langle s\rvert.$
@@
denote the probability of being at site $\lvert j\rangle$ at step $n,$ given that we started at site $\lvert i\rangle,$ with initial density $\rho$ concentrated at $\lvert i\rangle.$ Therefore, the dynamics starts with a density operator $\rho$ concentrated at some vertex $\lvert i\rangle,$ takes the evolution up to time ($n$) through the linear operator $\Phi^{n},$ producing a new density operator
@@
Consider an OQW on some set of vertices $\mathbb{Z}$, let $i\in\mathbb{Z}$ and $\rho\in\mathcal{D}(\mathfrak{h})$. We say that a vertex $\lvert i\rangle$ is
@@
    $\sum_{n=0}^{\infty}p_{ii;\rho}(n)=\infty.$ (2)
    Otherwise, $\lvert i\rangle$ is said to be $\rho$-transient;
@@
    recurrent, if $\lvert i\rangle$ is $\rho-$recurrent for all $\rho\in\mathcal{D}(\mathfrak{h});$
@@
    transient, if $\lvert i\rangle$ is $\rho-$transient for all $\rho\in\rho\in\mathcal{D}(\mathfrak{h}).$
Concerning all vertices, an OQW is called: recurrent if every vertex is recurrent; transient if every vertex is transient. Note that equation (2) computes the mean number of returns to site $\lvert i\rangle.$
The following Proposition describes $\rho$-recurrence in a concentrated vertex $i\in\mathbb{Z}$ when assumed that $\lvert i\rangle$ is recurrent for some density and $\mathfrak{h}$ has a finite internal degree of freedom. An immediate consequence will also be given for the case where $\lvert i\rangle$ is $\sigma$-transient for some non-faithful density.
@@
Consider an OQW on some set of vertices $V,$ let $i\in V$ and assume that $\mathfrak{h}$ is finite-dimensional. If $\lvert i\rangle$ is $\tau$-recurrent for some $\tau\in\mathcal{D}(\mathfrak{h}),$ then
@@
    vertex $\lvert i\rangle$ is $\rho$-recurrent for all faithful $\rho\in\mathcal{D}(\mathfrak{h})$;
@@
    there exists a pure density operator $\lvert w\rangle\langle w\rvert$ such that the OQW is $\lvert w\rangle\langle w\rvert$-recurrent, where $\lvert w\rangle$ is an unit vector of $\mathfrak{h}.$